# Knowledge Distillation on Spatial-Temporal Graph Convolutional Network for Traffic Prediction

### v2

@@ 3.2.2 Feature-based Knowledge Distillation in Hidden Layers
In these equations, $F\in\mathbb{R}^{B\times T\times N\times C}$, where $B$ is the batch size, $T$ is the number of time steps, $N$ is the number of graph nodes, and $C$ is the number of channels. Each element of the $TC$ tensor represents the absolute difference in features of a node at two time steps $i$ and $j$. The cost function $L_{\text{TCD}}$ addresses the difference in values for each node at two time steps. These errors are calculated using Equation 7 for both the teacher network and the student network. In the equation $\binom{T}{2}$, it accounts for the selection of any two time steps $i$ and $j$ out of $T$ due to normalization. Finally, the difference between these two sets of equations (8) yields the $L_{\text{TCD}}$ cost function.
Spatial Correlation Distillation Assuming a resemblance between the output of a spatial layer and the temporal layers, the cost function computes pairwise differences in values for all nodes at each time step in both the student and teacher networks (see Equation 9 ). These values are calculated for all spatial layers in both networks, and their pairwise differences are determined. To ensure a close alignment of the output from spatial layers in the student network with the corresponding layers in the teacher network, we introduce the cost function $L_{\text{SCD}}$, as defined in Equation 10 and illustrated in Figure 1.
@@
Each element of the $SC$ tensor represents the absolute difference in features of a node with another node at time step $t$. This tensor is computed for both the teacher and student networks (Equation 9). The $L_{\text{SCD}}$ cost function, calculated as the difference between these two vectors, is obtained from Equation 10.
Space-Time Cost Function Through the integration of three distillation components—overseen response distillation, temporal and spatial correlation distillation—we construct a comprehensive cost function aimed at distilling relevant information pertaining to relationships within both temporal and spatial layers. This facilitates the training of the student network to mimic the behavior of the teacher network. The derived cost function is defined in Equation 11 and depicted in Figure 1.
@@ 3.3 Pruning
In the knowledge distillation section, we introduced a cost function designed to train the student network by leveraging patterns identified by the teacher network. In this subsection, we present an algorithm that employs the $L_{\text{STCD}}$ cost function to train the student network. Additionally, it extracts the architecture of the student network by pruning insignificant parameters from the teacher network. The significance of each neuron is evaluated based on its influence on the network’s accuracy and its capacity to learn features extracted from the teacher. Considering a neural network with parameters $\mathbf{W}$ and a dataset $\mathcal{D}$, the training objective is to minimize the error $E$, as expressed in Equation 12:
[…]
In this equation, a regularization term is incorporated into the equation to reduce the number of parameters. The symbol $\lambda$ serves as a scaling factor, and $\|\cdot\|_{0}$ denotes the L0 norm, which signifies the count of non-zero elements. Minimizing the L0 norm poses a challenge due to its non-convex nature. As outlined in the method presented by [6], Equation 13 is utilized to establish a score $I$ for the parameter $m$. This involves employing gradients and weights obtained during the training phase through a first-order Taylor series expansion.
$\text{I}_{m}^{(1)}(W,B)=\sum_{b=1}^{B}(g_{m,b}\,w_{m,b})^{2}$ (13)
In this equation, $B$ represents the batch, $W$ is the weight matrix, and $g$ represents gradients. The importance score $\text{I}_{m}$ is straightforward to calculate since the gradient $g$ is obtained through backpropagation.
According to [6], this importance score can be thought of as a set of parameters labeled as $w_{s}$, similar to a convolutional filter. It is defined as a contribution to group sparsity in Equation 14:
@@
In this paper, Equation 14 is used to obtain the importance score of parameters. However, instead of using the gradients and weights obtained from training the network in the standard form, the network is trained using cost function defined in the section 3.2. The importance score obtained from the gradients and weights of the network, this time, not only indicates the importance of the parameter in the output but also reflects the learning ability of the knowledge perceived and extracted by the teacher. The proposed method is outlined in Algorithm 1.
@@
The algorithm defines a mask matrix $M$ to selectively retain or discard parameters in the network, with values of 1 or 0, respectively. An importance matrix KDIS is initialized to calculate the importance score of each parameter. The algorithm specifies minibatch intervals for pruning steps and the percentage of parameters to be pruned at each step. During each minibatch, the model undergoes fine-tuning using the $L_{\text{STCD}}$ loss function (line 11), where gradients and importance scores (KDIS) are computed based on the network weights. Knowledge distillation (KD) is applied to calculate gradients and weights, denoted as KDIS instead of I. After a specified minibatch count in line 5, importance values are averaged, and parameters are pruned in each layer according to the specified percentage (lines 13 to 19). The pruned network is then fine-tuned over multiple epochs (lines 24 and 25). The resulting mask matrix $M$ contains values of zero (removed parameters) and one (remaining parameters). Multiplying this matrix by the weight matrix of the student network’s base architecture yields the final pruned network.
@@ 4 Experiments and results
We thoroughly evaluate our model by testing it on two real traffic datasets: PeMSD7, which includes 228 nodes, and PeMSD8, with 170 nodes.
These datasets are collected by the California Department of Transportation and the San Bernardino City Transportation Commission.
PeMSD7 is collected by over 39,000 sensor stations located throughout the main urban areas of the California state freeway system [1]. This dataset is gathered at 5-minute intervals from 30-second data samples. In this paper, we randomly select an average scale from Region 7 of California, consisting of 228 stations, labeled as PeMSD7. The time range for the PeMSD7 dataset is on weekdays, during the months of May and June in the year 2012. We choose training, validation, and test sets of 35, 5, and 5 days, respectively.
PeMSD8 is similar to PeMSD7, with the difference that it includes traffic data for the city of San Bernardino from July to August 2016, collected from 170 monitoring stations on 8 streets at 5-minute intervals.
4.1 Data Preprocessing
A linear interpolation approach is employed to fill in missing values after data cleaning. Additionally, the input data is normalized using the Z-Score approach. The normalization process is represented in Equation 15:
$z=\frac{x-\mu}{\sigma}$ (15)
In this equation, $x$ represents the point for which we want to calculate its position relative to the dataset’s mean. $\mu$ denotes the mean of the dataset, serving as a reference point. $\sigma$ is the standard deviation of the dataset, representing a measure of the data’s dispersion. In PeMSD7 and PeMSD8, the adjacency matrix of the road graph is calculated based on the distances between stations in the traffic network. The weighted adjacency matrix $W$ can be formed by Equation 16:
$W_{ij}=\exp\left(-\frac{d_{ij}^{2}}{\sigma^{2}}\right)+\varepsilon$ (16)
Here, $W_{ij}$ is the weight of the edge determined by $d_{ij}$ (distance between station $i$ and $j$). $\sigma^{2}$ and $\varepsilon$ are thresholds controlling the distribution and dispersion of the $W$ matrix, assigned values of $10$ and $0.5$, respectively.
4.2 Hardware and Execution Details:
All presented results were generated using a TI 1080 GTX graphics card. We consistently used 12 historical timesteps $M$ to predict 9 future timesteps $H$. Our model predicts these 9 timesteps sequentially, as illustrated in Figure 2. In this figure, each prediction (e.g., $V_{t+1}$) is utilized as the last graph in the input series to forecast the next timestep ($V_{t+2}$). This sequential approach enables us to predict $h$ future timesteps. With data collected at 5-minute intervals, selecting 9 time units for prediction allows us to report results for the next 15, 30, and 45 minutes. Throughout all runs, the learning rate decreases by a factor of 0.7 after 5 epochs.
@@ 4.2 Hardware and Execution Details:
Visualization of Hidden Layer Representations All the results depicted in the t-SNE plots illustrate the hidden layer representation (which functions as the input to the output layer) for 50 test data points, each represented as an individual data instance, and is subsequently fed into t-SNE for visualization. In the plots, the red dots signify the student, while the blue dots represent either the teacher or the baseline pruning network (utilized exclusively for certain pruning results). The data employed in these visualizations is sourced from two datasets: PeMSD7 (depicted in Figure a) and PeMSD8 (depicted in Figure b). We can assess the effectiveness of our method by examining the t-SNE plots. In these visual representations, the proximity or overlap of the red and blue dots serves as an indicator of the model’s success. Ideally, when the method is effective, the red and blue dots should be close to or overlapping each other.
4.3 Evaluation Metrics and Training Details:
The results of traffic prediction in all tables are reported for the next 15, 30, and 45 minutes. The output values and calculated errors for these three time units are reported using three error functions: MAPE (Mean Absolute Percentage Error), MAE (Mean Absolute Error), and RMSE (Root Mean Square Error). The execution time reported in the tables are based on a batch of 1140 data and average over 100 runs.
4.4 Knowledge Distillation Analysis
All knowledge distillation experiments, except for pruning, employed the teacher and student models outlined in Table 1 and depicted in Figure 1 which provide comprehensive information on both networks, serving as a reference for subsequent comparisons and analysis. In the table, the column labeled ’Hidden Blocks Channel’ represents the channels for each spatial and temporal layer in the two hidden blocks of the ST-GCN network. For a more detailed explanation, we have visualized these channels on each layer in the accompanying Figure. Additionally, the columns: ”Test Time (s),” denotes the average test time over 100 trials, and ”FLOPS,” indicating the network’s floating-point operations per second (FLOPS) on a single forward pass. The student to teacher parameter ratio is approximately $3\%$ in both datasets. The execution time of the student network has reduced by approximately $83\%-84\%$ compared to the teacher. The number of blocks and layers in both networks is the same, and the difference in the number of parameters is due to the variation in the number of channels in each layer. The FLOPS ratio of the student network to the teacher is approximately $4\%$ for the PeMSD7 dataset and $2.5\%$ for the PeMSD8 dataset. We compared our proposed loss function, $L_{\text{ORD}}$, with other knowledge distillation loss functions such as $L_{\text{RD(KL)}}$, $L_{\text{RD(L2)}}$ [27, 28, 30, 29, 25]. The results are shown in Table 2. In all cases, except one (15-minute prediction on the PeMSD7 dataset), our approach has shown improvement compared to other approaches. For PeMSD7, $L_{\text{ORD}}$(ours) demonstrates competitive performance compared to $L_{\text{RD(L2)}}$ and $L_{\text{RD(KL)}}$ across all evaluated metrics. $L_{\text{ORD}}$(ours) consistently outperforming both $L_{\text{RD(L2)}}$ and $L_{\text{RD(KL)}}$ in MAE and RMSE for 15 min and 30 min predictions. This highlights the effectiveness of the overseen knowledge distillation method in capturing underlying patterns and improving accuracy. In the case of PeMSD8, $L_{\text{ORD}}$(ours) exhibits clear advantages over $L_{\text{RD(L2)}}$ and $L_{\text{RD(KL)}}$ in various metrics. Specifically, for MAE and RMSE at 15 min and 30 min prediction intervals, it consistently achieves lower errors, indicating superior predictive capabilities.
Table 2: Comparison of our approach $L_{\text{ORD}}$ with $L_{\text{RD(L2)}}$ and $L_{\text{RD(KL)}}$ on PeMSD7 and PeMSD8
   PeMSD7 Models MAPE MAE RMSE 15 min 30 min 45 min 15 min 30 min 45 min 15 min 30 min 45 min Teacher 5.223 7.316 8.739 2.230 3.010 3.565 4.097 5.752 6.834 Student W/O KD 6.423 9.685 12.298 2.666 3.868 4.799 4.649 6.938 8.610 Student $L_{\text{RD(L2)}}$ 6.379 9.661 12.474 2.768 4.214 5.527 4.709 7.185 9.178 Student $L_{\text{RD(KL)}}$ 6.411 9.527 11.894 2.700 3.938 4.918 4.672 6.984 8.657 Student $L_{\text{ORD}}$(ours) 6.411 9.516 12.104 2.762 4.091 5.221 4.645 6.847 8.569
   PeMSD8 Models MAPE MAE RMSE 15 min 30 min 45 min 15 min 30 min 45 min 15 min 30 min 45 min Teacher 2.293 3.239 3.925 1.211 1.665 2.031 2.524 3.501 4.081 Student W/O KD 2.967 4.035 4.734 1.472 1.956 2.294 2.988 4.090 4.706 Student $L_{\text{RD(L2)}}$ 2.812 3.976 4.908 1.426 1.953 2.387 2.839 3.978 4.733 Student $L_{\text{RD(KL)}}$ 2.964 4.179 5.057 1.507 2.125 2.575 2.895 3.971 4.664 Student $L_{\text{ORD}}$(ours) 2.661 3.717 4.553 1.363 1.885 2.285 2.788 3.862 4.573
We conducted a comparative analysis of our final loss function, $L_{\text{STCD}}$, against other approaches [27, 28, 30, 29, 25], and the results are presented in Table 4 and Figure 3. By reviewing the table, for PeMSD7, $L_{\text{STCD}}$(ours) demonstrates significant improvements over the baseline Student W/O KD and outperforms other knowledge distillation methods, including $L_{\text{RD(L2)}}$, $L_{\text{RD(KL)}}$, and SKD. Specifically, in terms of MAE and RMSE for the 15-minute and 30-minute predictions, $L_{\text{STCD}}$(ours) consistently achieves lower errors. Similarly, on the PeMSD8 dataset, $L_{\text{STCD}}$(ours) exhibits superior performance across various metrics when compared to $L_{\text{RD(L2)}}$, $L_{\text{RD(KL)}}$, and SKD. Noteworthy is the consistent outperformance in MAE and RMSE at the 15-minute and 30-minute prediction intervals, emphasizing the effectiveness of $L_{\text{STCD}}$(ours) in capturing spatial-temporal dependencies and enhancing predictive accuracy. These findings highlight the efficacy of the $L_{\text{STCD}}$(ours) approach, underscoring its potential as a robust spatial-temporal loss function for knowledge distillation in the context of traffic flow prediction datasets. Figure 3 illustrates better convergence and pattern imitation of the $L_{\text{STCD}}$ approach compared to other approaches in both knowledge bases. Additionally, we report the ratio of attention to teacher output instead of the dataset target in Table 3.
(a) PeMSD7
(b) PeMSD8
Figure 3: Comparison of our spatial-temporal correlation distillation loss function $L_{\text{STCD}}$ with the $L_{\text{RD(L2)}}$, $L_{\text{RD(KL)}}$ and $L_{\text{SKD}}$ loss functions.
Table 4: Performance comparison of our spatial-temporal correlation distillation loss function $L_{\text{STCD}}$ with other loss functions
   PeMSD7 Models MAPE MAE RMSE 15 min 30 min 45 min 15 min 30 min 45 min 15 min 30 min 45 min Teacher 5.223 7.316 8.739 2.230 3.010 3.565 4.097 5.752 6.834 Student W/O KD 6.423 9.685 12.298 2.666 3.868 4.799 4.649 6.938 8.610 Student $L_{\text{RD(L2)}}$ 6.379 9.661 12.474 2.768 4.214 5.527 4.709 7.185 9.178 Student $L_{\text{RD(KL)}}$ 6.411 9.527 11.894 2.700 3.938 4.918 4.672 6.984 8.657 Student $L_{\text{SKD}}$ 6.282 9.503 11.902 2.762 4.092 5.101 4.735 7.161 8.923 Student $L_{\text{STCD}}$(ours) 6.078 9.043 11.488 2.615 3.776 4.754 4.537 6.678 8.344
   PeMSD8 Models MAPE MAE RMSE 15 min 30 min 45 min 15 min 30 min 45 min 15 min 30 min 45 min Teacher 2.293 3.239 3.925 1.211 1.665 2.031 2.524 3.501 4.081 Student W/O KD 2.967 4.035 4.734 1.472 1.956 2.294 2.988 4.090 4.706 Student $L_{\text{RD(L2)}}$ 2.812 3.976 4.908 1.426 1.953 2.387 2.839 3.978 4.733 Student $L_{\text{RD(KL)}}$ 2.964 4.179 5.057 1.507 2.125 2.575 2.895 3.971 4.664 Student $L_{\text{SKD}}$ 2.726 3.819 4.553 1.409 1.944 2.317 2.909 4.139 4.894 Student $L_{\text{STCD}}$(ours) 2.491 3.375 4.067 1.281 1.719 2.052 2.716 3.707 4.385
Ablation Study Finally, we conducted an ablation study on the $L_{\text{STCD}}$ loss function, and the results are presented in Table 5. In PeMSD7, comparing various knowledge distillation models, $L_{\text{STCD}}$(ours) exhibits competitive performance against other variants, including $L_{\text{ORD}}$, $L_{\text{TCD}}$, and $L_{\text{SCD}}$. For instance, in terms of Mean Absolute Percentage Error (MAPE), Mean Absolute Error (MAE), and Root Mean Squared Error (RMSE) at different predictions (15 min, 30 min, and 45 min), $L_{\text{STCD}}$(ours) consistently demonstrates effectiveness in achieving lower errors compared to the baseline Student W/O KD and other knowledge distillation methods. Moving to the PeMSD8 dataset, the results in Table 5 further emphasize the robust performance of $L_{\text{STCD}}$(ours). When compared to $L_{\text{ORD}}$, $L_{\text{TCD}}$, and $L_{\text{SCD}}$, $L_{\text{STCD}}$(ours) consistently achieves lower MAPE, MAE, and RMSE across various predictions. Specifically, at 15 min and 30 min intervals, $L_{\text{STCD}}$(ours) outperforms its counterparts, showcasing its efficacy in capturing spatial-temporal dependencies and enhancing predictive accuracy.
Table 5: Results from the ablation study comparing the impact of individual loss functions ( $L_{\text{SCD}}$, $L_{\text{TCD}}$, and $L_{\text{ORD}}$ ), contributing to the construction of our overall $L_{\text{STCD}}$ loss function.
   PeMSD7 Models MAPE MAE RMSE 15 min 30 min 45 min 15 min 30 min 45 min 15 min 30 min 45 min Teacher 5.223 7.316 8.739 2.230 3.010 3.565 4.097 5.752 6.834 Student W/O KD 6.423 9.685 12.298 2.666 3.868 4.799 4.649 6.938 8.610 Student $L_{\text{ORD}}$ 6.411 9.516 12.104 2.762 4.091 5.221 4.645 6.847 8.569 Student $L_{\text{TCD}}$ 6.320 9.411 11.667 2.730 3.966 4.853 4.678 6.899 8.512 Student $L_{\text{SCD}}$ 6.326 9.193 11.380 2.743 3.957 4.866 4.645 6.853 8.476 Student $L_{\text{STCD}}$ 6.078 9.043 11.488 2.615 3.776 4.754 4.537 6.678 8.344
   PeMSD8 Models MAPE MAE RMSE 15 min 30 min 45 min 15 min 30 min 45 min 15 min 30 min 45 min Teacher 2.293 3.239 3.925 1.211 1.665 2.031 2.524 3.501 4.081 Student W/O KD 2.967 4.035 4.734 1.472 1.956 2.294 2.988 4.090 4.706 Student $L_{\text{ORD}}$ 2.661 3.717 4.553 1.363 1.885 2.285 2.788 3.862 4.573 Student $L_{\text{TCD}}$ 2.509 3.497 4.215 1.296 1.747 2.063 2.665 3.716 4.406 Student $L_{\text{SCD}}$ 2.619 3.680 4.457 1.355 1.850 2.214 2.722 3.778 4.478 Student $L_{\text{STCD}}$ 2.491 3.375 4.067 1.281 1.719 2.052 2.716 3.707 4.385
4.4.1 Hyperparameters
You can refer to Tables 6 for the learning rate, batch size, and hyperparameters related to knowledge distillation. These parameters are crucial in shaping the behavior of the models during training. In the context of the provided equations, the ”Models” column enumerates various knowledge distillation approaches evaluated for PeMSD7 and PeMSD8 datasets. Each approach, such as $L_{\text{RD(L2)}}$, $L_{\text{RD(KL)}}$, $L_{\text{SKD}}$, $L_{\text{ORD}}$(ours), and $L_{\text{STCD}}$(ours), incorporates distinct strategies for transferring knowledge. The ”Batch Size” column indicates the number of training samples utilized in each iteration for a specific model and dataset. For instance, SKD employs a batch size of 16 for PeMSD7 and 60 for PeMSD8. $L_{\text{RD(L2)}}$, $L_{\text{RD(KL)}}$, $L_{\text{ORD}}$(ours), and $L_{\text{STCD}}$(ours) maintain a batch size of 50 for both datasets, influencing the granularity of parameter updates during training. The ”Learning Rate” column specifies the step size during optimization, impacting the convergence of the models. Across various models and datasets, a learning rate of 1E-03 is consistently chosen, except for SKD on PeMSD8, which employs a learning rate of 1E-02. This parameter governs the size of the steps taken to reach a minimum in the loss landscape. The hyperparameters $\alpha_{1},\alpha_{2},\alpha_{3},\alpha,$ and $\beta$ are pivotal in the knowledge distillation process, as defined in equations (6), (11), (2), and (3). These values determine the contributions of different loss terms and the overall weighting of knowledge distillation components. For instance, $\alpha_{1}$ in $L_{\text{ORD}}$(ours) plays a role in deciding whether to prioritize teacher predictions or target labels based on a threshold.

### v4

@@ 3.1.2 Feature-based Knowledge Distillation in the Hidden Layers
In these equations, $F\in\mathbb{R}^{B\times T\times N\times C}$, where $B$ is the batch size, $T$ is the number of time steps, $N$ is the number of graph nodes, and $C$ is the number of channels. Each element of the $TC$ tensor represents the absolute difference in features of a node at two time steps $i$ and $j$. The cost function $L_{\text{TCD}}$ addresses the difference in values for each node at two time steps. These errors are calculated using Equation (7) for both the teacher network and the student network. In the equation $\binom{T}{2}$, it accounts for the selection of any two time steps $i$ and $j$ out of $T$ due to normalization. Finally, the difference between these two sets of Equations (8) yields the $L_{\text{TCD}}$ cost function.
Spatial Correlation Distillation Assuming a resemblance between the output of a spatial layer and the temporal layers, the cost function computes pairwise differences in values for all nodes at each time step in both the student and teacher networks (see Equation (9) ). These values are calculated for all spatial layers in both networks, and their pairwise differences are determined. To ensure a close alignment of the output from spatial layers in the student network with the corresponding layers in the teacher network, we introduce the cost function $L_{\text{SCD}}$, as defined in Equation (10) and illustrated in Figure 1.
@@
Each element of the $SC$ tensor represents the absolute difference in features of a node with another node at time step $t$. This tensor is computed for both the teacher and student networks (Equation (9)). The $L_{\text{SCD}}$ cost function, calculated as the difference between these two vectors, is obtained from Equation (10).
3.1.3 Space-Time Cost Function
Through the integration of three distillation components, we construct a comprehensive cost function aimed at distilling relevant information pertaining to relationships within both temporal and spatial layers. The derived cost function is defined in Equation (11) and depicted in Figure 1.
@@ 3.2 Pruning
In this subsection, we present an algorithm that employs the $L_{\text{STCD}}$ cost function to train the student network. Additionally, it extracts the architecture of the student network by pruning insignificant parameters from the teacher network. The significance of each neuron is evaluated based on its influence on the network’s accuracy and its capacity to learn features extracted from the teacher. As outlined in the method presented by Molchanov et al. [2019], Equation (12) is utilized to establish a score $I$ for a set of parameters labeled as $w_{s}$, similar to a convolutional filter. It is defined as a contribution to group sparsity in Equation (12):
@@
In this paper, Equation (12) is used to obtain the importance score of parameters. However, instead of using the gradients and weights obtained from training the network in the standard form, the network is trained using cost function defined in the section 3.2. The importance score obtained from the gradients and weights of the network, this time, not only indicates the importance of the parameter in the output but also reflects the learning ability of the knowledge perceived and extracted by the teacher. The proposed method is outlined in Algorithm 1. The algorithm defines a mask matrix $M$ to selectively retain or discard parameters in the network, with values of 1 or 0, respectively. An importance matrix KDIS is initialized to calculate the importance score of each parameter. The algorithm specifies minibatch intervals for pruning steps and the percentage of parameters to be pruned at each step. During each minibatch, the model undergoes fine-tuning using the $L_{\text{STCD}}$ loss function (line 11), where gradients and importance scores (KDIS) are computed based on the network weights. Knowledge distillation (KD) is applied to calculate gradients and weights, denoted as KDIS instead of I. After a specified minibatch count in line 5, importance values are averaged, and parameters are pruned in each layer according to the specified percentage (lines 13 to 19). The pruned network is then fine-tuned over multiple epochs (lines 24 and 25). The resulting mask matrix $M$ contains values of zero (removed parameters) and one (remaining parameters). Multiplying this matrix by the weight matrix of the student network’s base architecture yields the final pruned network.
@@ 4 Experiments and results
The results of traffic prediction in all tables are reported for the next 15, 30, and 45 minutes. The output values and calculated errors for these three time units are reported using three error functions: MAPE (Mean Absolute Percentage Error), MAE (Mean Absolute Error), and RMSE (Root Mean Square Error). The execution time reported in the tables are based on a batch of 1140 data and average over 100 runs. We thoroughly evaluate our model by testing it on two real traffic datasets: PeMSD7, which includes 228 nodes, and PeMSD8, with 170 nodes. PeMSD7 is collected by over 39,000 sensor stations located throughout the main urban areas of the California state freeway system Yu et al. [2018]. In this paper, we randomly select an average scale from Region 7 of California, consisting of 228 stations, labeled as PeMSD7. PeMSD8 is similar to PeMSD7, with the difference that it includes traffic data for the city of San Bernardino from July to August 2016, collected from 170 monitoring stations on 8 streets at 5-minute intervals. All presented results were generated using a TI 1080 GTX graphics card. We consistently used 12 historical timesteps $M$ to predict 9 future timesteps $H$. Our model predicts these 9 timesteps sequentially, as illustrated in Figure 2. In this figure, each prediction (e.g., $V_{t+1}$) is utilized as the last graph in the input series to forecast the next timestep ($V_{t+2}$). This sequential approach enables us to predict $h$ future timesteps. With data collected at 5-minute intervals, selecting 9 time units for prediction allows us to report results for the next 15, 30, and 45 minutes. Throughout all runs, the learning rate decreases by a factor of 0.7 after 5 epochs.
@@
4.1 Hyperparameters
You can refer to Tables 2 for the learning rate, batch size, and hyperparameters related to knowledge distillation. In the context of the provided equations, the ”Models” column enumerates various knowledge distillation approaches evaluated for PeMSD7 and PeMSD8 datasets. The hyperparameters $\alpha_{1},\alpha_{2},\alpha_{3},\alpha,$ and $\beta$ are pivotal in the knowledge distillation process, as defined in equations (6), (11), (2), and (3). The hyperparameters for the proposed pruning algorithm are presented in Table 3. The table categorizes hyperparameters based on the percentage of pruning applied, as well as batch size, learning rate, and three alpha values ($\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$), specific to the PeMSD7 and PeMSD8 datasets. For instance, in the PeMSD7 row with a pruning percentage of $97\%$, only $3\%$ of the model parameters are retained. The associated hyperparameters include a batch size of 25, a learning rate of $1\times 10^{-3}$, and alpha values $\alpha_{1}=0.099$, $\alpha_{2}=0.091$, and $\alpha_{3}=0.531$.